Imagine that you are visiting the planet Mars. You want to record your weight on this planet in a notebook  If your weight on earth is 450 N  What would you record as your weight on the Mars. You will find the required data and also take g=10m/s^2

Grade:11

1 Answers

Arun
5 years ago
Mass = 450 /10 = 45 kg
 
Weight on Mars = GMm /R²
 
On solving
 
 = 97.4 kg
 
Hope it helps
 
Regards
Arun (askIITianw forum expert)
 =